How old are Edz and Noe now, if Edz is twice as old as Noe and the sum of their ages is 66? ITS MY QUIZ. PLEASE ANSWER :)
Answer by Maths68(1474)

Let
Edz present age = e
Noe present age = n

Edz is twice as old as Noe
e=2n..............(1)
The sum of their ages is 66
e+n=66............(2)
Put the value of e in (2)
2n+n=66
3n=66
3n/3=66/3
n=22
Put the value of n in (1)
e=2n..............(1)
e=2(22)
e=44


Edz present age = e = 44
Noe present age = n = 22